Ÿ Households save 12% of their total income Ÿ Investment is given by� = 3500 − 10000�
Ÿ The tax rate is 20% Ÿ The real interest rate is 4% (r=0.4)
Ÿ Government expenditures are 2500
Ÿ Exports are 1900
Ÿ The import rate is 18%
a. Calculate the numerical value of the following and put your answer. You can show the
graph if you are able to draw using programs.
• Current level of investment
• Equilibrium aggregate demand
• Equilibrium personal saving
• The government budget surplus or deficit (indicate which) in equilibrium
• The change in equilibrium aggregate demand if the real interest rate rise to 8%
1
Expert's answer
2017-02-24T15:03:05-0500
a) Current level of investment
I=3500-10000r
I=3500-10000*0.04=3100

b) Equilibrium aggregate demand
S(personal)=I+NX-(T-G)
0.12Y=3100+1900-0.18Y-0.2Y+2500
Y=15000

c) Equilibrium personal saving
S(personal)=3100+1900-2700-3000+2500=1700

d) The government budget surplus in equilibrium
G-T=2500-0.2*15000=(-500)
G<T - budget surplus amounted to 500

e) The change in equilibrium aggregate demand if the real interest rate rise to 8%
I'=3500-10000*0.08=2700
0.2Y'=2700+1900-0.18Y'+2500-0.2Y'
Y'=14200
ΔY=14200-15000=-800 - equilibrium aggregate demand decreases by 800